Problem #1: How many years does it take for a deposit of $1000 to reach $3150 with an annual effective interest rate of 15%? Problem #1: 8.2097 Answer correct to 4 decimals Just Save Your work has been saved! (Back to Admin Page) Submit Problem #1 for Grading Problem #1 | Attempt #1 Your Answer: Attempt #2 | Attempt #3 | Attempt #4 | Attem pt #5 8.2096 8.2097 Your Mark: 0/2x 2/2 Problem #2: Redo Problem #1 above, compounding for each complete year, but using simple interest for any remaining fraction of a year at the end. (This is what is actually done in practice.) Problem #2 Answer correct to 4 decimals Just Save Your work has been saved! (Back to Admin Page) Submit Problem #2 for Grading Problem #2 | Attempt #1 Your Answer:14.3333 Attempt #2 8.2097 0/2x Attempt #3 | Attempt #4 | Attempt #5 21.0000 0/2x 1.0250 0/2x Your Mark: 0/2x

For problem #1 I used the formula FV=PV(I+i)^n where I solved for n to get 8.2097. Im unsure of how to solve question #2. Help ASAP would be very appreciated. Thank you.
Answer #1

As per question,

Target deposit = $3150

Interest Rate = 15%

with annual compounding upto 8 years,

FV = 1000(1.15)8

FV at the end of 8 years = $3059.02

As per question we need to use SI for fraction of a year,

So, required interest = 3150- 3059.02

Required interest = 90.98

As per Simple Interest Equation,

I = P * R * T

T = 90.98(0.15*1000)

T = 0.6065

So total time period = 8.6065 years